OpenStudy (anonymous):

how do I figure out the limit of n approaching infinity e^3/n+8

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ e ^{3} }{n+8 }\] One way to do this is to plug in infinity into n. Infinity + 8 is still essentially infinity. e^3 is a constant. Any constant divided by infinity is 0. (Imagine dividing a number by a very large number: the answer will be very small, and almost zero.) There are a few other ways but this is the simplest. Hope it helps!

OpenStudy (anonymous):

no it is e raise to the 3/n+8 that is what is throwing me off

OpenStudy (anonymous):

So the equation is: \[\lim_{n \rightarrow \infty}e ^{\frac{ 3 }{ n+8 }}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, we can still plug infinity into n, and using the same logic, we have: \[\lim_{n \rightarrow \infty} e ^{\frac{ 3 }{ \infty +8 }}\] \[\lim_{n \rightarrow \infty} e ^{\frac{ 3 }{ \infty }}\] \[\lim_{n \rightarrow \infty} e ^{0}\] Does this help?

OpenStudy (anonymous):

It's the same as \(\lim_{h\rightarrow0}e^h\) with \(h = \frac3{n+8}\).

OpenStudy (anonymous):

yes hold on...let me double check because I believe I tried that and webassign did not like the answer.

OpenStudy (anonymous):

thanks wotseit. I really dislike webassign. ugh. But that is correct...thanks

OpenStudy (anonymous):

No problem! Glad I could help :)

OpenStudy (phi):

the last step is change \[e^0 \] to a number for the answer.

OpenStudy (tkhunny):

You could also work with the logarithm of the expression.

OpenStudy (tkhunny):

It wouldn't hurt to remember your Order of Operations. You could have written: e^[3/(n+8)] Use parentheses to clarify intent.